Write electric field due to point charge.
1
Expert's answer
2020-09-17T14:29:57-0400

We know that according to Coulomb's law the electrostatic force between two charges q1 and q2 is

"F = k_e\dfrac{q_1q_2}{r^2}" , where r is the distance between two charges. If we take the direction of force into account, we get "\vec{F} = k_e \dfrac{q_1q_2}{r^3}\vec{r}" . Here ke is the Coulomb's constant (~8.99×109 N⋅m2⋅C−2).

The electric field of the charge q1is

"E = \dfrac{F}{q_2} = k_e\dfrac{q_1}{r^2}" . If q1 is positive, the direction of the field is "from q1", if it is negative, the direction is "towards q1".